OpenStudy (anonymous):

Complete the ordered pairs so that each is a solution of the given linear equation. The question is y=1/4x-3 and the points are (-8, ) , ( , 1)

OpenStudy (anwaarullah):

y=1/4x-3 as x=-8 y=-1/35 when x=-8 than y=-1/35 and when y=1 1=1/4x-3 4x-3=1 x=4 when x=1 y=4

OpenStudy (texaschic101):

y = 1/4x - 3 .....(-8, ?).....so sub in -8 for x y = 1/4(-8) - 3 y = -2 - 3 y = -5 so we have points (-8,-5) y = 1/4x - 3....(?, 1)....so sub in 1 for y 1 = 1/4x - 3 1 + 3 = 1/4x 4 = 1/4x 4/(1/4) = x 4 * 4/1 = x 16 = x so we have points (16,1) lets check.. slope(m) = (y2 - y1) / (x2 - x1) slope(m) = (1 - (-5) / (16 - (-8) slope(m) = (1 + 5) / (16 + 8) slope(m) = 6/24 slope(m) = 1/4 y = mx + b slope(m) = 1/4 use either of your points to find b (-8,-5)...x = -8 and y = -5 now sub -5 = 1/4(-8) + b -5 = -2 + b -5 + 2 = b -3 = b making your equation be : y = 1/4x - 3 (correct) answer is : (-8,-5) and (16,1)

OpenStudy (anwaarullah):

check,,,, when x=-8 than y=-1/35 -1/35=1/4x-3 xcos mutiply 4x-3=-35 4x=-32 x=-8 first check and like the second

OpenStudy (anwaarullah):

dear its 1/-8*4-3 -8*4=-32 and -32+-3=-35 so -1/35 (:>>

OpenStudy (texaschic101):

it is : y = (1/4)x - 3 with 1/4 being the slope so if you sub in -8 for x y = 1/4(-8) - 3 y = -8/4 - 3 y = -2 - 3 y = -5 and I checked my answer at the end at it comes out correctly.

OpenStudy (anwaarullah):

its 1/4x-3 not 1/4 *x-3

OpenStudy (anonymous):

Yes texaschic101 is correct as -5 is the correct answer and I double checked in the back of the book. I just didn't know how to solve it. I appreciate it! Thank you all :)

OpenStudy (anwaarullah):

yeah if its a slop formul form y=mx+b then 101 is right